# Dexterous Contact-Rich Manipulation via the Contact Trust Region

### v1

@@ I Introduction
At the heart of many iterative algorithms for manipulation lies the question: what is a good local description of contact mechanics for contact-rich manipulation that i) faithfully captures local behavior, and ii) is sufficiently simple for efficient and scalable computation? Many classical works have considered how contact forces, subject to friction cone constraints, can locally affect changes in manipuland configurations via the contact Jacobian [1, 2, 3]. These mechanics have been used heavily as local representations of contact for planning and control in subsequent works. For instance, classical grasping analysis and synthesis relies on local wrench-space arguments [4, 5]. More recently, many contact-implicit trajectory optimization formulations have considered both friction cone constraints and the unilateral nature of contact via complementarity constraints[6, 7, 8]. When iteratively solving the resulting optimization problems, the constraints are linearized, revealing similar contact Jacobians as their classical counterparts.
At surface level, these approaches seem different from those that use a differentiable simulator [9, 10, 11, 12, 13, 14, 15] to build Taylor approximations of the dynamics, thus hiding away the explicit consideration of friction cone, non-penetration and complementarity constraints from the planning algorithms. Previous works have utilized this locally-linear first-order Taylor approximation for the purposes of trajectory optimization and control [16, 17, 18, 11, 19, 20].
@@
The apparent discrepancy between these two approaches raises the question of how they are related. In fact, analytically deriving the gradients from differentiable simulators shows that the Taylor expansions they produce and the complementarity constraints ubiquitous in contact-implicit trajectory optimization are fundamentally connected[11, Example 5]. However, the form of each approximation hides a deeper question: where can we trust this local model? In optimization, this region is known as the trust region [21]. Intuitively, the trust region describes where the local model closely approximates the original function, making it reliable for local improvements. As the quality of Taylor approximations typically degrades as we move further from the nominal point, previous works have often employed an ellipsoidal trust region [21, 22, 23].
Our first contribution (Section III) is to elucidate the inconsistency between the ellipsoidal trust region and the unilateral nature of contact. We then address this issue by examining the structure of modern differentiable simulators and proposing the contact trust region (CTR). By augmenting Taylor expansions from these simulators explicitly with feasibility constraints from contact-implicit trajectory optimization, the CTR correctly characterizes the local region in which the Taylor expansions approximate the true contact dynamics well. Not only does CTR connect the Taylor expansions to classical concepts such as the wrench set, it can also be readily derived by combining standard components of existing differentiable simulators.
@@
In this work, we develop CTR based on the quasidynamic differentiable simulator proposed in [11], called the Convex Quasidynamic Differentiable Contact (CQDC) contact model. The approach could be extended to other contact dynamics models. The CQDC model, recapped in Section II, is very representative of approaches taken by modern differentiable simulators through contact [11, 12, 24, 25]. It treats simulation of contact as a convex optimization problem whose primal solution becomes the next state $x_{+}$, and gives gradients by performing sensitivity analysis in convex optimization.
@@
However, while a bulk of previous approaches [18, 11, 17, 19] have only focused on building a Taylor approximation of the configurations based on sensitivity gradients, our CTR also builds a linear model over the contact forces by utilizing sensitivity analysis to obtain gradients with respect to dual variables. A linear model over the contact forces can further qualify the ellipsoidal trust region by imposing friction cone constraints which are often present in works that directly encode contact complementarity constraints.
@@
The CTR and its connection to convex formulations of contact dynamics also allows us to seamlessly integrate the advances in contact smoothing for planning and control [12, 18, 11, 7, 26]. Without contact smoothing, many classical methods as well as local trajectory optimizers are limited to fixed contact-mode sequences. Thus, their ability to describe contact mechanics locally has been limited to the relationship between possible contact forces and changes in state within a contact mode. The myopicness of this local model is inconsistent with the non-smooth nature of contact, which also spells trouble for gradient-based Taylor approximations of contact dynamics.
@@
While complementarity-based representations [27] and mixed-integer formulations [28, 29, 30, 31] attempt to alleviate this limitation by considering contact mode changes globally, optimizing through these models often require searching through all possible contact-mode sequences. Given the near-exponential scalability of contact modes [32], these methods have struggled to tackle problems that are rich in contact. Thus, many successful algorithms go through an approximate representation in which complementarity is omitted in practice [33, 34, 35].
@@
To alleviate these limitations, smoothing-based methods relax contact dynamics by reasoning about the average result of all the contact modes around a nominal state given some distribution of state and action disturbances. Recent works [18, 9] have attributed the success and scalability of Reinforcement Learning (RL) to this mechanism. Surprisingly, however, [11, 36, 12] shows that for contact dynamics simulated by convex optimization, this process can be done without Monte-Carlo by using barrier smoothing which performs interior-point relaxation of the original contact dynamics and produces force from a distance. However, despite occasional successes, naively applying classical control methods such as LQR on smoothed linearization of contact dynamics often leads to catastrophic failures [37].
@@
As our second contribution, we present a highly efficient, contact-implicit Model Predictive Control (MPC) method [38, 33] in Section IV. Specialized for contact-rich manipulation, it can be viewed as a natural extension of iterative LQR [39]: we leverage Taylor expansions from CQDC dynamics as linear dynamics constraints, and we include the CTR as additional constraints to capture local contact dynamics. By formulating the CTR as a convex set—specifically, an intersection of multiple second-order cone constraints—each iteration of the proposed MPC remains a convex optimization problem.
We show the efficacy of the proposed CTR and MPC method on two representative contact-rich manipulation problems in Figure 1: i) whole-body manipulation on the bimanual iiwa and cylinder system, and ii) in-hand reorientation on the Allegro hand and cube system. For both systems, we aim to answer the following two questions: i) is the MPC method successful in planning for goals under quasidynamic equations of motion (Section V), and ii) can the MPC method successfully stabilize quasidynamic plans under true second-order dynamics (Section VI)? Through $2000$ trajectory runs in simulation and $100$ runs on hardware, our results delve into the efficacy and the limitations of our approach, as well as the benefits of CTR over the ellipsoidal one in the context of planning and control.
@@
As our final contribution, we address global planning by chaining local trajectories discovered by MPC. To do this, we follow a roadmap [40] approach where we seed each node with a stable object configuration. Then, we connect these nodes by first finding good actuator configurations for local MPC to start from using CTR (Section VII), then rolling out the MPC plan. To demonstrate the constructed roadmap’s robustness, we perform on the Allegro hand system $150$ consecutive edge traversals before the hardware shuts down due to overheating. For new goals that are not in the roadmap, we simply find the closest node to the goal, then connect them by the same process of finding a good initial actuator configuration and then rolling out MPC.
@@
Our proposed global planning method (Section VIII) enables efficient dexterous re-orientation on the Allegro hand, and has several advantages compared to existing approaches. Compared to single-query algorithms [41, 42, 11, 43], most of the computation time for our algorithm happens offline in the roadmap-building stage, enabling fast online inference. Moreover, the offline roadmap construction only takes minutes on a standard laptop, which is orders-of-magnitude less than approaches based on deep RL[44, 45, 46].
@@
Through our three contributions of i) the contact trust region that approximates contact mechanics efficiently, ii) a highly effective gradient-based MPC controller specialized to local contact-rich manipulation, and iii) a global planner that stitches together local trajectories, our method succeeds in dramatically reducing computation time compared to state-of-the-art methods in RL [47, 45, 46] or sampling-based MPC methods [48, 49] without the need for parallelization. We hope digging deep into the structure of contact mechanics can bring efficiency to the current state-of-the-art approaches for whole-body manipulation and dexterous in-hand reorientation.
II Contact Dynamics as Convex Optimization
@@
Many state-of-art methods for simulating contact solve a constrained optimization problem at every step [50, 51, 52, 10, 12, 53, 11]. In this paper, we use the quasi-dynamic formulation of contact dynamics presented in [11], where we assume the system is highly damped by frictional forces.
II-A Setup & Notation

### v3

@@ 1 Introduction
At the heart of many iterative algorithms for manipulation lies the question: what is a good local description of contact mechanics for contact-rich manipulation that (i) faithfully captures local behavior, and (ii) is sufficiently simple for efficient and scalable computation? Many classical works have considered how contact forces, subject to friction cone constraints, can locally affect changes in manipuland configurations via the contact Jacobian (Craig, 1986; Murray et al., 1994; Mason and Salisbury, 1985). These mechanics have been used heavily as local representations of contact for planning and control in subsequent works. For instance, classical grasping analysis and synthesis relies on local wrench-space arguments (Prattichizzo and Trinkle, 2008; Han et al., 2000). More recently, many contact-implicit trajectory optimization (CITO) formulations have considered both friction cone and the unilateral nature of contact via complementarity constraints (Ding et al., 2019; Posa et al., 2014; Wensing et al., 2024; Li et al., 2025). When iteratively solving the resulting optimization problems, the constraints are linearized, revealing similar contact Jacobians as their classical counterparts.
At surface level, these approaches seem different from those that use a differentiable simulator (Suh et al., 2022a; Pang and Tedrake, 2021; Pang et al., 2023; Howell et al., 2023a; Freeman et al., 2021; Macklin, 2022; Tedrake and the Drake Development Team, 2019) to build Taylor approximations of the dynamics, which replace the explicit consideration of friction cone, non-penetration and complementarity constraints with a simple linear equality constraint. Previous works have utilized this locally-linear first-order Taylor approximation for the purposes of trajectory optimization and control (Tassa et al., 2012; Howell et al., 2023b; Suh et al., 2022b; Pang et al., 2023; Kurtz and Lin, 2022; Kong et al., 2023).
@@
The apparent discrepancy between these two approaches raises the question of how they are related. In fact, analytically deriving the gradients from differentiable simulators shows that the Taylor expansions they produce and the complementarity constraints ubiquitous in contact-implicit trajectory optimization are fundamentally connected (Fiacco, 1983)(Pang et al., 2023, Example 5): the Taylor expansions are obtained by applying the Implicit Function Theorem to the force-balance and complementarity conditions, which form a subset of the contact dynamics constraints typically included in CITO formulations.
However, this connection hides a deeper question: where can we trust the Taylor expansions? In optimization, this region is known as the trust region (Sorensen, 1982). Intuitively, the trust region describes where the local model closely approximates the original function, making it reliable for local improvements. As the quality of Taylor approximations typically degrades as we move further from the nominal point, previous works have often employed an ellipsoidal trust region (ETR) (Sorensen, 1982; Nocedal and Wright, 2006; More, 1993).
Our first contribution (Section 3) is to reveal the inconsistency between a standard, symmetric ETR and the unilateral nature of contact. We address this by proposing the Contact Trust Region (CTR), which consists of a collection of convex constraints that correct this flaw. The key insight is that while the Implicit Function Theorem uses a subset of contact constraints to derive the Taylor expansion, the remaining feasibility constraints (e.g., friction) are merely ignored. The CTR re-incorporates these crucial, omitted constraints, augmenting the Taylor expansion to correctly characterize the local valid region. Not only does this connect the Taylor expansion to classical concepts like the wrench set, but it can also be readily derived from standard components of existing differentiable simulators. Furthermore, despite its convexity, we show that CTR is locally equivalent to the full set of smoothed contact dynamics constraint commonly used in CITO, allowing us to circumvent the numerically challenging complementarity constraints (Schwartz, 2011) when modeling contact.
To build intuition, consider a simple example of a spherical robot pushing a box across a table (see Figure 3 for a illustration of this system). At a state where the robot is touching the box, an ETR, due to its symmetric nature, would locally suggest a “pulling” action is as plausible as a “pushing” action. However, pulling on the object is physically impossible due to the nature of contact forces. The core idea of CTR is to build the physical constraint of contact directly into the trust region representation, pruning away physically infeasible actions like pulling.
@@
In this work, we develop CTR based on the quasidynamic differentiable simulator proposed in Pang et al. (2023), called the Convex Quasidynamic Differentiable Contact (CQDC) contact model. The CQDC model, recapped in Section 2, is very representative of approaches taken by modern differentiable simulators through contact (Pang et al., 2023; Howell et al., 2023a; Zhong et al., 2022; Werling et al., 2021), which makes extending CTR to full second-order dynamics conceptually straightforward. The CQDC model treats simulation of contact as a convex optimization problem whose primal solution becomes the next state $x_{+}$, and gives gradients by performing sensitivity analysis in convex optimization.
@@
However, while a bulk of previous approaches (Suh et al., 2022b; Pang et al., 2023; Howell et al., 2023b; Kurtz and Lin, 2022) have only focused on building a Taylor approximation of the configurations based on sensitivity gradients, our CTR also builds a linear model over the contact forces by utilizing sensitivity analysis to obtain gradients with respect to dual variables. A linear model over the contact forces can further qualify the ellipsoidal trust region by imposing friction cone constraints which are often present in works that directly encode contact complementarity constraints.
@@
The CTR and its connection to convex formulations of contact dynamics also allows us to seamlessly integrate the advances in contact smoothing for planning and control (Howell et al., 2023a; Suh et al., 2022b; Pang et al., 2023; Posa et al., 2014; Onol et al., 2019). Without contact smoothing, many classical methods as well as local trajectory optimizers are limited to fixed contact-mode sequences. Thus, their ability to describe contact mechanics locally has been limited to the relationship between possible contact forces and changes in state within a contact mode. The myopicness of this local model is inconsistent with the non-smooth nature of contact, which also spells trouble for gradient-based Taylor approximations of contact dynamics.
@@
While complementarity-based representations (Aydinoglu et al., 2022) and mixed-integer formulations (Marcucci et al., 2017; Marcucci and Tedrake, 2019; Marcucci et al., 2024; Graesdal et al., 2024) attempt to alleviate this limitation by considering contact mode changes globally, optimizing through these models often require searching through all possible contact-mode sequences. Given the near-exponential scalability of contact modes (Huang et al., 2020), these methods have struggled to tackle problems that are rich in contact. Thus, many successful algorithms go through an approximate representation in which complementarity is omitted in practice (Aydinoglu and Posa, 2022; Todorov, 2014; Jin, 2024; Kurtz et al., 2023).
@@
To alleviate these limitations, smoothing-based methods relax contact dynamics by reasoning about the average result of all the contact modes around a nominal state given some distribution of state and action disturbances. Recent works (Suh et al., 2022b, a) have attributed the success and scalability of Reinforcement Learning (RL) to this mechanism. Surprisingly, however, Pang et al. (2023); Zhang et al. (2023); Howell et al. (2023a) show that for contact dynamics simulated by convex optimization, this process can be done without Monte-Carlo by using barrier smoothing which performs interior-point relaxation of the original contact dynamics and produces force from a distance. However, despite occasional successes, naively applying classical control methods such as LQR on smoothed linearization of contact dynamics often leads to catastrophic failures (Shirai et al., 2024a).
@@
As our second contribution, we present a highly efficient, contact-implicit Model Predictive Control (MPC) method (Le Cleac’h et al., 2024; Aydinoglu and Posa, 2022) in Section 4. Specialized for contact-rich manipulation, it can be viewed as a natural extension of iterative LQR (Li and Todorov, 2004): we leverage Taylor expansions from CQDC dynamics as linear dynamics constraints, and we include the CTR as additional constraints to capture local contact dynamics. By formulating the CTR as a convex set—specifically, an intersection of multiple second-order cone constraints—each iteration of the proposed MPC remains a convex optimization problem.
We show the efficacy of the proposed CTR and MPC method on two representative contact-rich manipulation problems in Figure 1: i) whole-body manipulation on the bimanual iiwa and cylinder system, and ii) in-hand reorientation on the Allegro hand and cube system. For both systems, we aim to answer the following two questions: i) is the MPC method successful in planning for goals under quasidynamic equations of motion (Section 5), and ii) can the MPC method successfully stabilize quasidynamic plans under true second-order dynamics (Section 6)? Through $2000$ trajectory runs in simulation and $100$ runs on hardware, our results delve into the efficacy and the limitations of our approach, as well as the benefits of CTR over the ellipsoidal one in the context of planning and control.
@@
As our final contribution, we address global planning by chaining local trajectories discovered by MPC. To do this, we follow a roadmap (Kavraki et al., 1996) approach where we seed each node with a stable object configuration. Then, we connect these nodes by first finding good actuator configurations for local MPC to start from using CTR (Section 7), then rolling out the MPC plan. To demonstrate the constructed roadmap’s robustness, we perform on the Allegro hand system $150$ consecutive edge traversals before the hardware shuts down due to overheating. For new goals that are not in the roadmap, we simply find the closest node to the goal, then connect them by the same process of finding a good initial actuator configuration and then rolling out MPC.
@@
Our proposed global planning method (Section 8) enables efficient dexterous re-orientation on the Allegro hand, and has several advantages compared to existing approaches. Compared to single-query algorithms (Cheng et al., 2022, 2024; Pang et al., 2023; Khandate et al., 2024), most of the computation time for our algorithm happens offline in the roadmap-building stage, enabling fast online inference. Moreover, the offline roadmap construction only takes minutes on a standard laptop, which is orders-of-magnitude less than approaches based on deep RL(OpenAI et al., 2019; Chen et al., 2022; Handa et al., 2022).
@@
Through our three contributions of (i) the contact trust region that approximates contact mechanics efficiently (Section 2,3), (ii) a highly effective gradient-based MPC controller specialized to local contact-rich manipulation (Section 4-6), and (iii) a global planner that stitches together local trajectories (Section 7,8), our method succeeds in dramatically reducing computation time compared to state-of-the-art methods in RL (OpenAI et al., 2019; Chen et al., 2022; Handa et al., 2022) or sampling-based MPC methods (Howell et al., 2022; Li et al., 2024) without the need for parallelization. We hope digging deep into the structure of contact mechanics can bring efficiency to the current state-of-the-art approaches for whole-body manipulation and dexterous in-hand reorientation.
Table 1: List of Symbols
Symbol Shape Meaning
$n_{q}$ $\mathbb{N}$ number of positions in a system
$n_{q^{\mathrm{a}}}$ $\mathbb{N}$ number of actuated positions
$n_{q^{\mathrm{o}}}$ $\mathbb{N}$ number of object positions
$q$ $\mathbb{R}^{n_{q}}$ system configuration vector, $q\coloneqq({q^{\mathrm{a}}},{q^{\mathrm{o}}})$
$q^{\mathrm{a}}$} $\mathbb{R}^{n_{q_{\mathrm{a}}}}$ actuated (robot) configuration vector
$q^{\mathrm{o}}$} $\mathbb{R}^{n_{q_{\mathrm{o}}}}$ object configuration vector
$q_{+}^{\mathrm{o}/\mathrm{a}}$ $\mathbb{R}^{n_{q_{\mathrm{o}}}/n_{q_{\mathrm{a}}}}$ object/robot configuration at the next step
$u$ $\mathbb{R}^{n_{q_{\mathrm{a}}}}$ control input
$\lambda_{i}$ $\mathbb{R}^{3}$ contact force at contact $i$, $\lambda_{i}\coloneqq(\lambda_{\mathrm{n}_{i}},\lambda_{\mathrm{t}_{i}})$
$\lambda_{\mathrm{n}_{i}}$ $\mathbb{R}$ normal contact force at contact $i$
$\lambda_{\mathrm{t}_{i}}$ $\mathbb{R}^{2}$ tangential contact force at contact $i$
$\phi_{i}$ $\mathbb{R}$ signed distance at contact $i$, (5a)
$\tau^{\mathrm{o}/\mathrm{a}}$ $\mathbb{R}^{n_{q_{\mathrm{o}}}/n_{q_{\mathrm{a}}}}$ non-contact torque acting on object/robot
$h$ $\mathbb{R}_{+}$ step size in seconds
$\mathbf{K}_{\mathrm{a}}$ $\mathbf{S}^{n_{q_{\mathrm{a}}}}_{++}$ robot stiffness matrix
$\mathbf{M}_{\mathrm{o}}$ $\mathbf{S}^{n_{q_{\mathrm{o}}}}_{++}$ object mass matrix
$\mathbf{J}_{i}$ $\mathbb{R}^{3\times n_{q}}$ contact Jacobian at contact $i$, (2)
$\mathcal{K}_{i}$ $\mathbb{R}^{3}$ feasible velocity cone at contact $i$, (4a)
$\mathcal{K}_{i}^{\star}$ $\mathbb{R}^{3}$ friction cone at contact $i$, (4b)
$\mu_{i}$ $\mathbb{R}_{+}$ friction coefficient at contact $i$
$\nu_{i}$ $\mathbb{R}^{3}$ (5a)
$\mathbf{P}$ $\mathbf{S}^{n_{q}}_{++}$ quadratic cost in contact dynamics, (3b)
$b$ $n_{q}$ linear cost in contact dynamics, (3c)
$c_{i}$ $\mathbb{R}^{3}$ (7c)
$\kappa$ $\mathbb{R}_{+}$ contact dynamics smoothing parameter
$f_{(\kappa)}(\cdot,\cdot)$ $\mathbb{R}^{n_{q}}\times\mathbb{R}^{n_{q_{\mathrm{a}}}}$ ($\kappa$-smoothed) contact dynamics, ((14))(8)
$\rightarrow\mathbb{R}^{n_{q}}$
$\mathbf{A}_{(\kappa)}$ $\mathbb{R}^{n_{q}\times n_{q}}$ $\partial f_{(\kappa)}/\partial q$
$\mathbf{B}_{(\kappa)}$ $\mathbb{R}^{n_{q}\times n_{q_{\mathrm{a}}}}$ $\partial f_{(\kappa)}/\partial u$
$\mathbf{C}_{(\kappa),i}$ $\mathbb{R}^{n_{q}\times n_{q}}$ $\partial\lambda_{(\kappa),i}/\partial q$
$\mathbf{D}_{(\kappa),i}$ $\mathbb{R}^{n_{q}\times n_{q_{\mathrm{a}}}}$ $\partial\lambda_{(\kappa),i}/\partial u$
$\bar{\cdot}$ N.A. nominal quantities, e.g. $\bar{q}$, see (17)
$\delta{\cdot}$ N.A. deviation from nominal quantities, e.g. $\delta{q}$
$\hat{\cdot}$ N.A. Taylor expansions, e.g. $\hat{q}_{+}$
$\mathbf{\Sigma}$ $\mathbf{S}_{++}^{(n_{q}+n_{q_{\mathrm{a}}})}$ trust region radius
$\mathcal{E}_{\mathbf{\Sigma}}$ $\mathbb{R}^{n_{q}+n_{q_{\mathrm{a}}}}$ ETR with radius $\mathbf{\Sigma}$, (22)
$\mathcal{S}_{\mathbf{\Sigma},\kappa}$ $\mathbb{R}^{n_{q}+n_{q_{\mathrm{a}}}}$ CTR with radius $\mathbf{\Sigma}$ and smoothing $\kappa$, (24)
$\mathcal{S}_{\mathbf{\Sigma},\kappa}^{\mathcal{A}}$ $\mathbb{R}^{n_{q_{\mathrm{a}}}}$ Action-only CTR, (25)
$\tilde{\mathcal{S}}_{\mathbf{\Sigma},\kappa}$ $\mathbb{R}^{n_{q}+n_{q_{\mathrm{a}}}}$ Relaxed CTR, (26)
$\tilde{\mathcal{S}}_{\mathbf{\Sigma},\kappa}^{\mathcal{A}}$ $\mathbb{R}^{n_{q_{\mathrm{a}}}}$ Relaxed Action-only CTR, (27)
$\mathcal{M}_{\mathbf{\Sigma,\kappa}}^{(\mathcal{A})}$ $\mathbb{R}^{n_{q}}$ (action-only) motion set, ((29))(28)
$\mathcal{C}^{\mathcal{A}}_{\mathbf{\Sigma},\kappa,i}$ $\mathbb{R}^{3}$ contact force set at contact $i$, (30)
$\mathcal{JC}^{\mathcal{A}}_{\mathbf{\Sigma},\kappa,i}$ $\mathbb{R}^{n_{q_{\mathrm{o}}}}$ generalized friction cone at contact $i$, (31)
$\mathcal{W}^{\mathcal{A}}_{\mathbf{\Sigma},\kappa}$ $\mathbb{R}^{n_{q_{\mathrm{o}}}}$ wrench set, (32)
$\mathcal{M}^{\mathcal{A},\mathrm{o}}_{\mathbf{\Sigma},\kappa}$ $\mathbb{R}^{n_{q_{\mathrm{o}}}}$ motion set projected to the space of $q^{\mathrm{o}}$}, (33)
$T$ $\mathbb{N}$ CtrTrajOpt planning horizon in Algorithm 1
$n_{\mathrm{max}}$ $\mathbb{N}$ CtrTrajOpt iteration limit in Algorithm 1
$H$ $\mathbb{N}$ MPC rollout horizon in Algorithm 2
$r$ $\mathbb{R}_{+}$ scalar trust region radius in Section 5.2.1
$\mathbf{Q}$ $\mathbf{S}_{+}^{n_{q}}$ state tracking quadratic in (37)
$\mathbf{R}$ $\mathbf{S}_{+}^{n_{q_{\mathrm{a}}}}$ control action quadratic cost in (37)
NN $\mathbb{N}$ number of MPC re-plans in Algorithm 3
2 Preliminaries: Contact Dynamics as Convex Optimization
@@
Many state-of-art methods for simulating contact solve a constrained optimization problem at every step (Stewart and Trinkle, 2000; Anitescu, 2006; Todorov et al., 2012; Pang and Tedrake, 2021; Howell et al., 2023a; Han et al., 2023; Pang et al., 2023). In this paper, we use the quasi-dynamic formulation of contact dynamics presented in (Pang et al., 2023), where we assume the system is highly damped by frictional forces.
2.1 Setup & Notation